Solve: 5(1 - 5x) + 5(-8x + 2) = -4x – 8x. Round your answer to four decimal places
Answer by MichaelaButz(2)

5(1 - 5x) + 5(-8x + 2) = -4x – 8x
When you distribute this out you get...
5-25x-40x+10=-4x-8x
Then you combine all like terms to get...
(5+10)-(25x-40x)=-12x
15-65x=-12x
Now the next goal is to get the variable x all by itself on one side of the equation...
15=53x
(15/53)=x
Then plug that into your calculator and round it to four decimals :).

Hope that helped!